has stock outstanding as follows: 25,000 shares of cumulative preferred 2% stock, $150 par, and 83,000 shares of $10 par common. During its first four years of operations, the following amounts were distributed as $109,750; third year, $155,630; fourth year, $216,100 Compute the dividends per share on each class of stock for each of year, enter o the four years. Round all answers to two dedimal places. If no dividends are paid in a given 1st Year 2nd Year 3rd Year 4th Year
Answer #1
Shares Outstanding
Cumulative preferred 2% stock,$150 par----25,000shares
Total preferred share capital=150*25,000=$3,750,000
Common shares, $10 par----83,000shares
FIRST YEAR
Dividend distributed in first year=$50,250
Dividend due on preferred stock=2% ie,( 0.02*$3,750,000)=$75,000
Hence entire $50,250 distributed to preferred stockholder
Dividend per share on preferred stock=$50250/25,000=$2.01
Dividend on common stock=Nil
Dividend Preferred Common
Per share Stock Stock
First year $2.01 Nil
SECOND YEAR
Arrear of dividend on preferred share=75000-50250=$24,750
This years dividend due=$75000
Total dividend due on cumulative preferred shares=$75,000+$24,750=$99750
Actual dividend paid during the year=$109750
Dividend paid to preferred shareholders=$99750
Dividend paid to common shareholders=$109750-99750= $10000
Dividend per share paid to preferred shareholder=$99750/25,000=$3.99
Dividend per share paid to common shareholder=$10000/83000 = $0.12
Dividend Preferred Common
Per share Stock Stock
Second Year $3.99 $0.12
THIRD YEAR
Dividend due for preferred shareholders=$75000
Dividend distributed =$155630
Dividend paid to common shareholders=(155630-75000)=$80630
Dividend per share paid to preferred shareholder=75000/25000 = $3.000
Dividend per share paid to common shareholder=$80630/83000 = $0.97
Dividend Preferred Common
Per share Stock Stock
Third Year $3.00 $0.97
FOURTH YEAR
Dividend due for preferred shareholders=$75000
Dividend distributed =$216100
Dividend paid to common shareholders=216100-75000 141100
Dividend per share paid to preferred shareholder=75000/25000 = $3.00
Dividend per share paid to common shareholder=$141100/83000 = $1.70
Dividend Preferred Common
Per share Stock Stock
Fourth Year $3.00 $1.70
YEAR WISE DIVIDEND PER SHARE
Dividend Preferred Common
Per share Stock Stock
First year $2.01 Nil
Second Year $3.99 $0.12
Third Year $3.00 $0.97
Fourth Year $3.00 $1.70